# Scalable learning of potentials to predict time-dependent Hartree-Fock dynamics

### v1

@@ II.3 Generation of data sets for training and testing
To train and test our models, we follow procedures for data generation, taking care to keep these procedures uniform across all molecular systems that we study. In what follows, $H(P)$ can either be the ground truth Hamiltonian, or it can be one of the $\widetilde{H}(P)$ models described above.
@@
Ensemble of field-free trajectories: We take the $P(0)$ described in the previous paragraph and perturb it using a random Hermitian matrix $\mathscr{H}$, creating a new initial condition $\mathscr{P}(0)=P(0)+\varepsilon\mathscr{H}$. To generate $\mathscr{H}$, we start with $2N^{2}$ independent samples from the standard normal distribution, from which we form two real matrices $\mathscr{D}_{R}$ and $\mathscr{D}_{I}$ both of size $N\times N$. Let denote conjugate transpose. We take $\mathscr{D}=\mathscr{D}_{R}+\imath\mathscr{D}_{I}$ and then set $\mathscr{H}=(\mathscr{D}+\mathscr{D}^{\dagger})/2$. In this paper, we set $\varepsilon=(10/N^{2})\sum_{i,j}|P_{ij}(0)|$, i.e., $10$ times the mean absolute value of $P(0)$. Though $\mathscr{P}(0)$ is Hermitian, we must modify it slightly to be idempotent, a requirement for Hartree-Fock electron density matrices. We compute the eigendecomposition $\mathscr{P}(0)=VDV^{\dagger}$. Any entries of $D$ greater than $1/2$ are set to $1$, while entries at most equal to $1/2$ are set to 00. With this modified $D$, the resulting $\mathscr{P}(0)=VDV^{\dagger}$ is guaranteed to be idempotent.
@@
When we numerically solve (2) in any setting, we monitor $P(t)$ at every time step to ensure that it remains Hermitian and idempotent.
II.4 Training
@@
To train a given $\widetilde{H}(P;\bm{\theta})$ model, we need training data. For a given molecular system (choice of molecule plus basis set), we use the true Hamiltonian (8) in the field-free procedures described in Section II.3. This yields two different training sets (single field-free trajectory and ensemble of field-free trajectories). Both training sets are of the form $\{(P^{(j)},\dot{P}^{(j)})\}_{j=1}^{n_{\text{train}}}$; each pair in this set consists of an electron density matrix and its time-derivative, taken from the same trajectory and evaluated at the same time.
@@ II.4.1 Gradient- and Hessian-based minimization of the loss
As $\mathbf{Hess}$ is of size $|\bm{\theta}|\times|\bm{\theta}|$, we can only use (25a) to train our models when the number of parameters $|\bm{\theta}|$ is sufficiently small. When $\mathbf{Hess}$ has eigenvalues near zero, we replace the matrix inverse with the Moore-Penrose pseudoinverse (Hansen, 1998) with threshold $10^{-12}$. Here we treat problems for which $N$ can be as large as $60$; in this case, for the models (11) and (16), $|\bm{\theta}|=N^{4}=12.96\times 10^{6}$. Assuming $8$-byte double precision floats, a Hessian of size $N^{4}\times N^{4}$ would consume over $1$ petabyte, motivating the use of matrix-free methods.
@@ II.4.2 Matrix-free minimization of the loss
Returning to the loss (24), we see that the critical objects are the Jacobians $A^{j}=\nabla_{\bm{\theta}}S^{(j)}$, which can be written as $A^{j}_{k,l,m}={\partial S^{(j)}_{kl}}/{\partial\theta_{m}}$. Thus $\mathbf{A}$ is a $4$-index tensor of dimension $n_{\text{train}}\times N\times N\times|\bm{\theta}|$. Since we will not (in general) be able to store $\mathbf{A}$ in memory, we will present efficient numerical algorithms to compute the left and right products of $\mathbf{A}$ against tensors of appropriate sizes. We will use these algorithms as inputs to the matrix-free method LSMR (Fong and Saunders, 2011), and thereby compute minimizers of the loss (24).
Note that in situations where $\mathbf{A}$ does not have full column rank, LSMR returns the solution $\bm{\theta}^{\star}$ with minimal $2$-norm; this is particularly useful for problems such as ours, where $\mathbf{Hess}$ may have nearly zero eigenvalues and high condition number. Our prior work combatted these problems by combining Hessian-based training with a Tikhonov or ridge regularization approach (Gupta et al., 2022). For the present work, when we tried adding such regularization terms to the loss, we found that results did not improve, likely due to the implicit regularization already present in LSMR.
Please see SI Section V.3 for detailed expressions of Jacobian-vector products (both left and right) for all three models (11), (16), and (19). In SI, we also explain how we implement these products efficiently, so that we can train our models even when they contain tens of millions of parameters. For model (19), we also provide (in SI Section V.3.3) expressions for the gradient and Hessian of the loss $\mathcal{L}$, which are used to train model (19) for small molecular systems.
III Results
We test our models on seven molecular systems, each consisting of a molecule together with a particular atom-centered Gaussian basis set: (i) $\text{HeH}^{+}$ with 6-31G ($N=4$), (ii) $\text{HeH}^{+}$ with 6-311G ($N=14$), (iii) LiH with 6-31G ($N=11$), (iv) $\text{C}_{2}\text{H}_{4}$ with STO-3G ($N=14$), (v) LiH with 6-311++$\text{G}^{**}$ ($N=29$), (vi) $\text{C}_{2}\text{H}_{4}$ with 6-31+$\text{G}^{*}$ ($N=46$) and (vii) $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$ with STO-3G ($N=60$). In SI Section V.5, we provide the Cartesian coordinates for each molecule.
III.1 Model training
@@
The ensemble training set initially consists of a cloud of $100$ matched trajectories $\{\mathscr{P}_{k}(t_{j}),\dot{\mathscr{P}}_{k}(t_{j})\}_{k=1}^{100}$ for $j=2,\ldots,J-2$ with $J=20000$. For the smaller molecular systems, we take every $50$th snapshot from this cloud, for a total of $40000$ temporal snapshots ($400$ snapshots from each of the $100$ trajectories). We augment this using $40000$ snapshots from the single field-free trajectory and its time-derivative, i.e., we augment using every $5$th snapshot from $\{P(t_{j}),\dot{P}(t_{j})\}_{j=2}^{j=199998}$. In total, for smaller molecular systems, our ensemble training set consists of $80000$ snapshots. For the larger molecular systems, we take every $100$ snapshot from the cloud and augment with every $10$th snapshot from the single field-free trajectory, leading to a total ensemble training set size of $40000$ snapshots.
@@
III.2 Metrics
In Figure 2, for the models trained with the iterative LSMR method, we plot training loss as a function of the number of iterations. Across all molecular systems, the SymmE model achieves the best results. For larger molecular systems, the SymmE model requires orders of magnitude fewer number of iterations to achieve small values of the training loss. For smaller molecular systems, any of the models trained on the ensemble of field-free trajectories (TiedE, HermE, and SymmE) yield the best performance in terms of rapid minimization of training loss, i.e., within $10^{3}$ iterations. For a detailed table and plot of final values of the training loss for all models and all molecular systems, see SI Section V.4.1.
After training, we assess model errors using three metrics: propagation error, Hamiltonian error, and commutator error.
(a) Field-free propagation errors
(b) Field-on propagation errors
III.2.1 Propagation Error
$\imath\frac{d\widetilde{P}(t)}{dt}=\big{[}\widetilde{H}(t,\widetilde{P}(t)),\widetilde{P}(t)\big{]},$ (26)
with a Hamiltonian $\widetilde{H}(t,P)$ that equals the learned field-free Hamiltonian $\widetilde{H}(P)$ plus an optional time-dependent term. As explained in Section II, we use the CI4 scheme to solve both (2) and (26) for $J=20000$ time steps at $\Delta t=8.268\times 10^{-4}$, starting both trajectories at the same initial condition $\widetilde{P}(0)=P(0)$. We then compute the $\infty$-norm error
the worst-case error across all matrix entries and all time.
In Figure 3, we show that in terms of propagation error (27) on both the field-free/training set (left) and the field-on/test set (right), the LSMR-trained models do no worse—and in some cases outperform—the Hessian-trained models. These results justify our use of the LSMR method to train the larger molecular systems, for which a comparison against the Hessian-based approach is impossible. For the remainder of this paper, we focus our attention on LSMR-trained models. Further results for Hessian-trained models can be found in SI Section V.4.2.
(a) Field-free propagation errors
(b) Field-on propagation errors
In Figure 4, we plot the propagation error (27) for all LSMR-trained models and all seven molecular systems, on both the field-free/training set (left) and the field-on/test set (right). All reported errors are orders of magnitude lower than what we achieved in prior results (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022). In nearly all cases, and especially for the three largest systems, ensemble-trained models (TiedE, HermE, and SymmE) outperform single trajectory-trained models (Tied, Herm, and Symm), often by several orders of magnitude. For the two small systems $\text{HeH}^{+}$ and LiH in 6-31G, the only model that outperforms ensemble-trained models is Symm, the $8$-fold symmetry-preserving model trained on a single field-free trajectory.
In SI Section V.4.2, we plot—for each model and each molecular system—the time-dependent mean absolute error (MAE)
III.2.2 Hamiltonian and Commutator Errors
For each of the three models we have proposed, there exists a set of parameters $\bm{\theta}^{\circ}$ such that the model perfectly matches the ground-truth Hamiltonian (8)—for detailed derivations of this, see SI Section V.2. Thus it makes sense to quantify, for each model, how far the trained parameter values $\bm{\theta}^{\star}$ are from the ground truth parameters $\bm{\theta}^{\circ}$. For the three respective models (11), (16), and (19), we compute the $\infty$-norm error
For each model, we view the parameters as a flattened vector $\bm{\theta}$ of length $|\bm{\theta}|$. We use $\bm{\theta}^{\ast}$ and $\bm{\theta}^{\circ}$ to denote, respectively, the trained and true values of the parameters.
In the left-most plot in Figure 5, we find that SymmE, the $8$-fold symmetry-preserving model (19) trained on the ensemble of field-free trajectories, has the smallest Hamiltonian error (29) across all molecular systems and all LSMR-trained models.
Finally, note that the field-free and field-on commutator errors show extremely similar trends. This is because, in the field-on case, the electric field terms in $H$ and $\widetilde{H}$ cancel perfectly, so that the commutator error reduces to $\|[H(P(t))-\widetilde{H}(P(t)),P(t)]\|$ where $H$ and $\widetilde{H}$ are the field-free Hamiltonians and $P(t)$ is the field-on trajectory.
Further tables of results can be found in SI Section V.4.3.
@@ IV Discussion and Conclusion
In our prior work, we always trained our models on single field-free trajectories (Bhat, Ranka, and Isborn, 2020; Gupta et al., 2022). The results presented in Section III show the importance of training with an ensemble of field-free trajectories. In Section II.3, we described the construction of this ensemble in detail. In particular, when we constructed random initial conditions for the cloud of trajectories in the ensemble, we ensured that these initial conditions were Hermitian and idempotent. Indeed, if we train on ensembles generated from trajectories that ignore these constraints, the resulting models perform poorly. Especially for larger molecular systems, training on physically meaningful ensembles constrains our models to stay closer to the true Hamiltonian than is possible with a single trajectory training set.
@@
We have focused on a method (TDHF) in which the underlying symmetries are known exactly. In cases where we seek to learn a term (e.g., an exchange-correlation potential or self-energy) whose full symmetries are unknown, the present paper offers two approaches. First, if there is a symmetry that one suspects should be present, one can easily generalize our methods to create an appropriately symmetry-constrained machine learning model. For a linear model, the only substantive changes to Algorithm 1 will concern the dimensions of the tensors and the elements of the symmetry group $\{\sigma^{m}\}$. For nonlinear neural network models, one can use or adapt methods explored previously (Anderson, Hy, and Kondor, 2019; Pan and Kondor, 2022). If the resulting model yields results that are significantly better than non-symmetrized models, then a reasonable working hypothesis is that the unknown term truly possesses the enforced symmetry.
Second, our results show that preserving only the Hermitian symmetry leads to predictive accuracy in terms of long-term propagation error: the TiedE and HermE models perform well in propagation error and commutator error tests. Though they cannot compare with SymmE in terms of ease of training (training loss versus iteration number) or $\infty$-norm Hamiltonian error, the TiedE and HermE models were far easier to implement (in terms of code development time) than the SymmE model. Generalizing to more complex situations such as TDDFT or TDRDMFT, we hypothesize that a machine learning model that preserves only an outer layer of symmetries may be sufficient for a first round of results. In subsequent rounds, these models should be refined by incorporating more symmetries, especially if the goal is to learn the true Hamiltonian. This will ensure that the model can extrapolate beyond the training data, and also increase the chances that one can distill new physical insights from, e.g., a machine-learned closure model for theories such as TDDFT and TDRDMFT.

### v2

@@ II.3 Generation of data sets for training and testing
To train and test our models, we generate data using the following procedures, taking care to keep these procedures uniform across all molecular systems that we study. At a high level, we emphasize that we train models using field-free trajectories, either a single such trajectory or an ensemble. We then test our models using field-on trajectories.
In what follows, $H(P)$ can either be the ground truth Hamiltonian, or it can be one of the $\widetilde{H}(P)$ models described above. When we numerically solve (2) in any setting, we monitor $P(t)$ at every time step to ensure that it remains Hermitian and idempotent, a requirement for Hartree-Fock electron density matrices.
@@
Ensemble of field-free trajectories: We take the $P(0)$ described in the previous paragraph and perturb it using a random Hermitian matrix $\mathscr{H}$, creating a new initial condition $\mathscr{P}(0)=P(0)+\varepsilon\mathscr{H}$. To generate $\mathscr{H}$, we start with $2N^{2}$ independent samples from the standard normal distribution, from which we form two real matrices $\mathscr{D}_{R}$ and $\mathscr{D}_{I}$ both of size $N\times N$. Let denote conjugate transpose. We take $\mathscr{D}=\mathscr{D}_{R}+\imath\mathscr{D}_{I}$ and then set $\mathscr{H}=(\mathscr{D}+\mathscr{D}^{\dagger})/2$. In this paper, we set $\varepsilon=(10/N^{2})\sum_{i,j}|P_{ij}(0)|$, i.e., $10$ times the mean absolute value of $P(0)$. Though $\mathscr{P}(0)$ is Hermitian, we must modify it slightly to be idempotent. We compute the eigendecomposition $\mathscr{P}(0)=VDV^{\dagger}$. Any entries of $D$ greater than $1/2$ are set to $1$, while entries at most equal to $1/2$ are set to 00. With this modified $D$, the resulting $\mathscr{P}(0)=VDV^{\dagger}$ is guaranteed to be idempotent.
@@
II.4 Training: Mathematical Formulation
@@
To train a given $\widetilde{H}(P;\bm{\theta})$ model for a given molecular system (choice of molecule plus basis set), we use one of two training sets: the single field-free trajectory or the ensemble of field-free trajectories. As detailed above in Section II.3, we generated both training sets by solving (2) with the true Hamiltonian (8). Both training sets are of the form $\{(P^{(j)},\dot{P}^{(j)})\}_{j=1}^{n_{\text{train}}}$; each pair in this set consists of an electron density matrix and its time-derivative, taken from the same trajectory and evaluated at the same time.
@@ II.4.1 Gradient- and Hessian-based minimization of the loss
As $\mathbf{Hess}$ is of size $|\bm{\theta}|\times|\bm{\theta}|$, we can only use (25a) to train our models when the number of parameters $|\bm{\theta}|$ is sufficiently small. When $\mathbf{Hess}$ has eigenvalues near zero, we replace the matrix inverse with the Moore-Penrose pseudoinverseHansen (1998) with threshold $10^{-12}$. Here we treat problems for which $N$ can be as large as $60$; in this case, for the models (11) and (16), $|\bm{\theta}|=N^{4}=12.96\times 10^{6}$. Assuming $8$-byte double precision floats, a Hessian of size $N^{4}\times N^{4}$ would consume over $1$ petabyte, motivating the use of matrix-free methods.
@@ II.4.2 Matrix-free minimization of the loss
Returning to the loss (24), we see that the critical objects are the Jacobians $A^{j}=\nabla_{\bm{\theta}}S^{(j)}$, which can be written as $A^{j}_{k,l,m}={\partial S^{(j)}_{kl}}/{\partial\theta_{m}}$. Thus $\mathbf{A}$ is a $4$-index tensor of dimension $n_{\text{train}}\times N\times N\times|\bm{\theta}|$. Since we will not (in general) be able to store $\mathbf{A}$ in memory, we will present efficient numerical algorithms to compute the left and right products of $\mathbf{A}$ against tensors of appropriate sizes. We will use these algorithms as inputs to the matrix-free method LSMR Fong and Saunders (2011), and thereby compute minimizers of the loss (24).
Note that in situations where $\mathbf{A}$ does not have full column rank, LSMR returns the solution $\bm{\theta}^{\star}$ with minimal $2$-norm; this is particularly useful for problems such as ours, where $\mathbf{Hess}$ may have nearly zero eigenvalues and high condition number. Our prior work combatted these problems by combining Hessian-based training with a Tikhonov or ridge regularization approach Gupta et al. (2022). For the present work, when we tried adding such regularization terms to the loss, we found that results did not improve, likely due to the implicit regularization already present in LSMR.
Please see SM Section VI.3 for detailed expressions of Jacobian-vector products (both left and right) for all three models (11), (16), and (19). In SM, we also explain how we implement these products efficiently, so that we can train our models even when they contain tens of millions of parameters. For model (19), we also provide (in SM Section VI.3.3) expressions for the gradient and Hessian of the loss $\mathcal{L}$, which are used to train model (19) for small molecular systems.
(a) Field-free propagation errors
(b) Field-on propagation errors
II.5 Training: Computational Workflow
@@
The ensemble training set initially consists of a cloud of $100$ matched trajectories $\{\mathscr{P}_{k}(t_{j}),\dot{\mathscr{P}}_{k}(t_{j})\}_{k=1}^{100}$ for $j=2,\ldots,J-2$ with $J=20000$. For the smaller molecular systems, we take every $50$th snapshot from this cloud, for a total of $40000$ temporal snapshots ($400$ snapshots from each of the $100$ trajectories). We augment this using $40000$ snapshots from the single field-free trajectory and its time-derivative, i.e., we augment using every $5$th snapshot from $\{P(t_{j}),\dot{P}(t_{j})\}_{j=2}^{j=199998}$. In total, for smaller molecular systems, our ensemble training set consists of $80000$ snapshots. For the larger molecular systems, we take every $100$th snapshot from the cloud and augment with every $10$th snapshot from the single field-free trajectory, leading to a total ensemble training set size of $40000$ snapshots.
@@
III Results
We test our models on seven molecular systems, each consisting of a molecule together with a particular atom-centered Gaussian basis set: (i) $\text{HeH}^{+}$ with 6-31G ($N=4$), (ii) $\text{HeH}^{+}$ with 6-311G ($N=14$), (iii) LiH with 6-31G ($N=11$), (iv) $\text{C}_{2}\text{H}_{4}$ with STO-3G ($N=14$), (v) LiH with 6-311++$\text{G}^{**}$ ($N=29$), (vi) $\text{C}_{2}\text{H}_{4}$ with 6-31+$\text{G}^{*}$ ($N=46$) and (vii) $\text{C}_{6}\text{H}_{6}\text{N}_{2}\text{O}_{2}$ with STO-3G ($N=60$). In SM Section VI.5, we provide the Cartesian coordinates for each molecule.
(a) Field-free propagation errors
(b) Field-on propagation errors
III.1 Metrics
For detailed tables/plots of final values of the training loss for all models and all molecular systems, see SM Section VI.4.1. After training, we assess model errors using three metrics: propagation error, Hamiltonian error, and commutator error.
III.1.1 Propagation Error
$\imath\frac{d\widetilde{P}(t)}{dt}=\big{[}\widetilde{H}(\widetilde{P}(t),t),\widetilde{P}(t)\big{]},$ (26)
with a Hamiltonian $\widetilde{H}(P,t)$ that equals the learned field-free Hamiltonian $\widetilde{H}(P)$ plus an optional time-dependent term. As explained in Section II, we use the CI4 scheme to solve both (2) and (26) for $J=20000$ time steps at $\Delta t=8.268\times 10^{-4}$, starting both trajectories at the same initial condition $\widetilde{P}(0)=P(0)$. We then compute the $\infty$-norm error
the worst-case error across all matrix entries and all time.
In Figure 2, we show that in terms of propagation error (27) on both the field-free/training set (left) and the field-on/test set (right), the LSMR-trained models do no worse—and in some cases outperform—the Hessian-trained models. These results justify our use of the LSMR method to train the larger molecular systems, for which a comparison against the Hessian-based approach is impossible. For the remainder of this paper, we focus our attention on LSMR-trained models. Further results for Hessian-trained models can be found in SM Section VI.4.2.
In Figure 3, we plot the propagation error (27) for all LSMR-trained models and all seven molecular systems, on both the field-free/training set (left) and the field-on/test set (right). All reported errors are orders of magnitude lower than what we achieved in prior results Bhat, Ranka, and Isborn (2020); Gupta et al. (2022). In nearly all cases, and especially for the three largest systems, ensemble-trained models (TiedE, HermE, and SymmE) outperform single trajectory-trained models (Tied, Herm, and Symm), often by several orders of magnitude. For the two small systems $\text{HeH}^{+}$ and LiH in 6-31G, the only model that outperforms ensemble-trained models is Symm, the $8$-fold symmetry-preserving model trained on a single field-free trajectory.
In SM Section VI.4.2, we plot—for each model and each molecular system—the mean absolute errors (MAE)
III.1.2 Hamiltonian and Commutator Errors
For each of the three models (11), (16), and (19), there exist parameters $\bm{\theta}^{\circ}$ such that the model perfectly matches the ground-truth Hamiltonian (8)—for detailed derivations, see SM Section VI.2. Hence the $\infty$-norm error between trained ($\bm{\theta}^{\star}$) and true ($\bm{\theta}^{\circ}$) parameters is one way to measure the error between trained and true Hamiltonians:
In the left-most plot in Figure 4, we find that SymmE, the $8$-fold symmetry-preserving model (19) trained on the ensemble of field-free trajectories, has the smallest Hamiltonian error (29) across all systems and all LSMR-trained models.
The field-free and field-on commutator errors show extremely similar trends. This is because, in the field-on case, the electric field terms in $H$ and $\widetilde{H}$ cancel perfectly, so that the commutator error reduces to $\|[H(P(t))-\widetilde{H}(P(t)),P(t)]\|$ where $H$ and $\widetilde{H}$ are the field-free Hamiltonians and $P(t)$ is the field-on trajectory.
@@ III.1.2 Hamiltonian and Commutator Errors
Further tables of Hamiltonian and commutator error results can be found in SM Section VI.4.3.
III.1.3 Iteration Count
In Figure 5, for the models trained with the iterative LSMR method, we plot training loss as a function of the number of iterations. Across all molecular systems, the SymmE model achieves the best results. For larger molecular systems, the SymmE model requires orders of magnitude fewer number of iterations to achieve small values of the training loss. For smaller molecular systems, any of the models trained on the ensemble of field-free trajectories (TiedE, HermE, and SymmE) yield the best performance in terms of rapid minimization of training loss, i.e., within $10^{3}$ iterations.
@@ IV Discussion and Conclusion
In our prior work, we always trained our models on single field-free trajectories Bhat, Ranka, and Isborn (2020); Gupta et al. (2022). The results presented in Section III show the importance of training with an ensemble of field-free trajectories. In Section II.3, we described the construction of this ensemble in detail. In particular, when we constructed random initial conditions for the cloud of trajectories in the ensemble, we ensured that these initial conditions were Hermitian and idempotent. Indeed, if we train on ensembles generated from trajectories that ignore these constraints, the resulting models perform poorly. Especially for larger molecular systems, training on physically meaningful ensembles constrains our models to stay closer to the true Hamiltonian than is possible with a single trajectory training set.
@@
We have focused on a method (TDHF) in which the underlying symmetries are known exactly. In cases where we seek to learn a term (e.g., an exchange-correlation potential or self-energy) whose full symmetries are unknown, the present paper offers two approaches that complement prior methods to learn symmetries from dataLiu and Tegmark (2022); Desai, Nachman, and Thaler (2022). First, if there is a symmetry that one suspects should be present, one can easily generalize our methods to create an appropriately symmetry-constrained machine learning model. The only substantive changes to Algorithm 1 will concern the dimensions of the tensors and the elements of the symmetry group $\{\sigma^{m}\}$. For nonlinear neural networks, our techniques can be incorporated in a linear output layer. One can then test empirically whether the resulting model yields results that are significantly better than non-symmetrized models.
Second, our results show that preserving only the Hermitian symmetry leads to predictive accuracy in terms of long-term propagation error: the TiedE and HermE models perform well in propagation error and commutator error tests. Though they cannot compare with SymmE in terms of computational cost (training loss versus iteration number) or $\infty$-norm Hamiltonian error, the TiedE and HermE models required less code development time than the SymmE model. Generalizing to more complex situations such as TDDFT or TDRDMFT, we hypothesize that a machine learning model that preserves only an outer layer of symmetries may be sufficient for a first round of results. In subsequent rounds, these models can be refined by incorporating more symmetries, especially if the goal is to learn the true Hamiltonian.
Both of these approaches, along with careful attention to generating an accurate ensemble of trajectories, will ensure that the model can extrapolate beyond the training data, and also increase the chances that one can distill new physical insights from, e.g., a machine-learned closure model for theories such as TDDFT and TDRDMFT. The scalability of our approach could be improved by reducing memory requirements with tensor train techniques or by density matrix dimensionality reduction, allowing application to larger molecular systems. However, as our goal is to use this method to derive new insights into how to learn an unknown molecular potential, such as the time-dependence of the exchange-correlation potential in TDDFT, molecules of the size explored here provide a rich playing field for exploration and study for future work.